a truck carries a 5.2 ton load of gravel.with this load it accelerates at 0.36m/s without a load the truck accelerate at 0.62m/s.what is the mass of the truck?the accelerating total force are the same in both area


1
Expert's answer
2022-02-05T10:27:37-0500

Newton's second law is the key:


"a_1=\dfrac F{m+M},\\\space\\
a_2=\dfrac FM,\\\space\\
\dfrac {a_1}{a_2}=\dfrac{M}{m+M}→M=\dfrac{a_1m}{a_2-a_1}=7.2\text{ t}."